priyankaa invested $2000, part at 8% per annum and the rest at 10% per annum. After one year the total interest earned was $190. How much did she invest at each rate?
Answer by stanbon(75887)

priyankaa invested $2000, part at 8% per annum and the rest at 10% per annum. After one year the total interest earned was $190. How much did she invest at each rate?
---------
Quantity Eq: x + y = 2000
Interest Eq:8x +10y = 19000
------------------------------
Multiply thru the Quantity Eq by 8:
8x + 8y = 16000
8x +10y = 19000
----
2y = 3000
y = $1500 (amt. of 10% investment)
----
solve for "x":
x + y = 2000
x + 1500 = 2000
x = $500 (amt. of 8% investment)
===========================
Cheers,
Stan H.
==============================